# RobMOT: Robust 3D Multi-Object Tracking by Observational Noise and State Estimation Drift Mitigation on LiDAR PointCloud

### v3

@@ 2.2 Refined Kalman filter for Noise Mitigation
The permitted observations are associated with trajectories cached in the framework. We use the Hungarian algorithm [18] with the algebraic association strategy followed in [17]. The association for an observation is performed by selecting a trajectory with the shortest Euclidean distance between the estimated state and the observation. When the shortest distance to an observation exceeds $\sigma$, a new trajectory is created for the observation.
The framework performs next-state estimation for all trajectories cached in the system. Trajectories associated with an observation will pass through state update and state prediction as demonstrated in Figure 2. Other trajectories will go directly through state prediction. The framework utilizes KF with a constant acceleration motion model.
We have observed drift in state estimation for objects encountering occlusion or leaving the sensor field of view (off-scene). This phenomenon is also observed in the literature, as displayed in Figure 1 when the estimated trajectory of a parked car with ID 4 drifts away from the expected location during the occlusion. Figure 3 shows another scenario when two parked cars with ID 0 and 2 drift from the actual position after being off-scene. Our investigation of the phenomenon concludes that this deviation is caused by trembling in detection bounding boxes around objects’ locations, which results in illusion movement for these objects that eventually impacts state update and prediction in KF.
We model this noise using the algebraic deviation, shown in Equation 1, between the detected bounding boxes $z^{det}_{ij}=(x_{ij},y_{ij})$ and the actual location of the objects $z^{gt}_{ij}=(x_{ij},y_{ij})$ across $k$ observations for $N$ objects using KITTI training dataset.
@@
As depicted in Figure 4, the error distribution shows a Gaussian shape for the noise of the VirConv detector across $x$ and $y$ LiDAR coordinates, which facilitates its integration into KF. Accordingly, we derive the $2\times 2$ covariance matrix $D_{t}$ representing the noise impact on the detections (measurements). Notably, the noise varies from one detector to another as to the variance in detection performance; hence, $D_{t}$ should be recomputed as the employed detector changes. By applying Equation 1 on $y$ coordinate, we obtain a covariance matrix of the noise, Equation 2.
@@
Since this noise is attached to the incoming detections $z_{t}$ at teim $t$, it harms the computation of the residual term $\hat{y}_{t}$, which is the difference between the current measurement $z_{t}$ and state estimation $\hat{x}_{t|t-1}$ predicted at $t-1$. Consequently, the noise impact on the residual term leads to incorrect updates for the motion parameters of the object. The innovation covariance $S_{t}$ is responsible for weighting the innovation residual $\hat{y}_{t}$ through the Kalman gain $K_{t}$. Therefore, the covariance matrix $D_{t}$ needs to be added to the innovation covariance $S_{t}$ to compensate for the impact of noise on the updated state. Hence, the updated state estimation $\hat{x}_{t|t}$ will consider the impact of the deviation noise on the observation.
@@
The main reason for not using $R_{t}$ instead is because it models the incorrectness of the measurement for the point cloud from LiDAR with respect to the real world, while $D_{t}$ models the incorrectness of the detection process with respect to the LiDAR point cloud, which means they are independent. Equation 3 demonstrates the updated state of KF after adapting $D_{t}$ noise.
$H_{t}$ is the observation model to map from estimation to measurement space at time $t$, while $P_{t|t-1}$ is the estimation uncertainty at time $t$ given measurement at $t-1$. In Section 2.4, we will describe how $P_{t|t}$ contribute to trajectory termination in the cache.
2.3 Trajectory Validity
The objective is to monitor the validity (legitimacy) of the existing trajectories in the system while identifying invalid ones, called ghost tracks. Each track has a certainty score that represents its legitimacy. This score is updated whenever a new observation of the track is available. The certainty score changes based on the likelihood of the trajectory being a ghost track. By observing the characteristics of ghost tracks, the ghost tracks are usually formed from intermittent sequences of observations with low to moderate detection scores. Based on these characteristics, the certainty score for track $i$ changes when a new observation is associated with the track at time $t$. The certainty score changes according to the detection score $s_{i}^{t}$ and the duration of absence $d^{i}_{t}$ since the last observation associated with the track.
@@ 2.3.2 Absence duration
The penalty term $\frac{d^{i}_{t}}{s^{i}_{t}}$ lowers the certainty score as the absence duration $d^{i}_{t}$ of a new observation for the track increases. This term will diminish the impact of the reward with an extended absence duration that indicates a high potential of being a ghost trajectory. However, this term can harm legitimate tracks that encounter occlusion. Since observations of legitimate tracks tend to have high detection scores $s^{i}_{t}$, we include inverse proportional to the detection score to decrease the impact of this term on the legitimate tracks.
@@ 2.3.3 Temporal validity
A visual representation of the distribution of Equation 4 is highlighted in Figure 5; the relation between the detection score, the absence duration, and its impact on the certainty score.
This mechanism monitors the validity of trajectories in the system individually until a certainty score of a trajectory exceeds a pre-defined threshold $\alpha_{legit}$ that indicates the trajectory is legitimate. In this case, the validation mechanism changes the trajectory status from “not confirmed” to “confirmed” and terminates the validation mechanism to this trajectory. Thus, confirmed trajectories cannot be changed to “not confirmed" as the mechanism is terminated for these tracks. Algorithm 2 demonstrates trajectory validation procedures in detail.
Algorithm 2 Trajectory Validity
  function UpdateTrajectoryCertainty($T,\alpha_{legit}$)
  Input:
   $T$: List of trajectories in the cache
   $\alpha_{legit}$: Trajectory legitimacy threshold
  for each $traj$ in $T$ do
     if $traj$.hasNewObservation() then
        $s_{t}\leftarrow traj$.getObservation().detectionScore()
        $d_{t}\leftarrow traj$.getAbsenceDuration()
        $f_{k}\leftarrow traj$.getCertaintyScore()
        $f_{t}=s_{t}e^{-d_{t}}-\frac{d_{t}}{s_{t}}+f_{k}$
        $traj$.setCertaintyScore($f_{t}$)
        if $traj$.getCertaintyScore() $\leq\alpha_{legit}$ then
           $traj$.setValidityStatus(true)
        end if
     end if
  end for
  end function
@@ 2.4 Trajectory Termination and Caching
All trajectories are saved in a cache module. In the literature, a trajectory is terminated or discarded when there are no further observations for that trajectory for several frames. Generally, it is not trivial to define an absence threshold as it could risk eliminating objects under prolonged occlusion or maintaining ghost tracks for a long. Instead, in our framework, we discard the trajectory when the estimation uncertainty, $P_{t|t}$ in Equation 3, becomes high, as it can lead to mismatching in object association. We have observed that ghost tracks tend to have limited observations, leading to high uncertainty in state estimation that matches the intended goal. Meanwhile, legitimate tracks maintain low estimation uncertainty, even those under occlusion. Thus, estimation uncertainty is used to terminate tracks from the cache. Algorithm 3 demonstrates the termination procedures in the cache.
Algorithm 3 Trajectory Termination
  function trajectory Termination($T,\alpha_{cov}$)
  Input:
   $T$: List of trajectories in the cache
   $\alpha_{cov}$: Estimation uncertainty threshold to terminate a trajectory
  for each $traj$ in $T$ do
     $estCertaintyX\leftarrow traj.$getCovariance_P(0)
     $estCertaintyY\leftarrow traj.$getCovariance_P(1)
     $discard\leftarrow$ false
     if $estCertaintyX>\alpha_{conv}$ then
        $discard\leftarrow$ true
     end if
     if $estCertaintyY>\alpha_{conv}$ then
        $discard\leftarrow$ true
     end if
     if $discard$ then
        Discard $traj$ from $T$
     end if
  end for
  end function
@@ 3 Results and Discussion
The framework is evaluated using two benchmark datasets for autonomous cars: KITTI [24] and Waymo Open Dataset [25] (WOD). The KITTI dataset consists of $50$ sequences, split into $21$ training sequences and $29$ testing sequences recorded around Karlsruhe, Germany, that involve a variety of challenging conditions. Meanwhile, WOD contains about $1150$ segments recorded in various locations across the United States. The experiments are conducted on a machine with a processor AMD Ryzen 9; no GPU is involved since the framework only utilizes the CPU.
@@ 3.1.1 Quantitative Evaluation
Table 1: The table compares the latest benchmarks listed in the KITTI leaderboard on the testing split. The highest performance at each metric is highlighted in red, the second highest in blue, and the third highest in green. Methods that use the CasA [20] detector are marked by (*), while ($\dot{+}$) is used for methods that use the VirConv [19] detector.
@@
Method HOTA MOTA AssA AssRe IDSW TripletTrack [26] 73.58% 84.32% 74.66% 77.3% 322 PolarMOT [27] 75.2% 85.1% 76.95% 80.0% 462 DeepFusion-MOT [2] 75.5% 84.6% 80.1% 82.6% 84 Mono-3D-KF [28] 75.5% 88.5% 77.6% 80.2% 162 PC3T [29] 77.8% 88.8% 81.6% 84.8% 225 MSA-MOT [30] 78.5% 88.0% 82.6% 85.2% 91 UG3DMOT* [31] 78.6% 87.98% 82.28% 85.36% 30 CasTrack* [1] 80% 90.5% 82.7% 86.4% 150 Rethink MOT [5] 80.1% 91.5% 83.6% 87.6% 46 MCTrack [32] 80.78% 89.82% 84.29% - 65 PC-TCNN [33] 80.9% 91.7% 84.1% 87.5% 37 $\dot{+}$ [19, 1] 81.56% 90.54% 85.2% 88.7% 113 RobMOT* (Ours) 81.22% 90.5% 85.8% 89.7% 7 $\dot{+}$ (Ours) 81.76% 91.02% 85.6% 89.25% 7
We evaluate the performance of the proposed framework, RobMOT [nagy2024robmotrobust3dmultiobject], against several state-of-the-art multi-object tracking benchmarks on the KITTI dataset testing split. The results are summarized in Table 1, which presents metrics: HOTA (Higher Order Tracking Accuracy), MOTA (Multi-Object Tracking Accuracy), AssA (Association Accuracy), AssRe (Association Recall), and IDSW (Identity Switches).
In this comparison, we use two different detectors, VirConv [19] and CasA [20], to evaluate the tracking performance rather than the detection fairly. We are re-running the official implementation of the CasTrack [1] work as it is the latest top-performing classical approach using their proposed detectors CasA [20] and VirConv [19]. Our proposed framework, RobMOT, outperforms CasTrack with the CasA detector, marked by (*) in Table 1, with about 1.22% in HOTA. By switching the detector to VirConv, our method still outperforms CasTrack [1], named "VirConvTrack" [19] in Table 1, with about 0.2% in HOTA, marked by ($\dot{+}$) in Table 1. These results show the contribution of the proposed tracking approach regardless of the performance of the utilized detector. Notably, RobMOT is ranked the first or second highest in most metrics, with CasA or VirConv detector among the other benchmarks.In addition, the proposed work excels in Association Accuracy (AssA) with 85.8%, marking it the highest among the evaluated methods. Similarly, it achieves a leading Association Recall (AssRe) of 89.7%, indicating its robust capability to maintain correct identity associations throughout the tracking process.
One of the most significant contribution is the low Identity Switches (IDSW) score of 7 for RobMOT with CasA and VirConv detectors. This demonstrates the model’s ability to consistently track object identities over time, which is critical for reliable multi-object tracking. In contrast, other benchmark methods, such as UG3DMOT [31] and CasTrack [1] with the same detector, exhibit substantially higher IDSW scores of 30 and 150, respectively, indicating more frequent identity switches.
Our further experiments on the KITTI evaluation dataset across five detectors show CasTrack [1] and VirConvTrack [19] dependency on the utilized detector, as illustrated in Figure 6. It shows the tracking performance across the detectors using three different metrics: MOTA, HOTA, and IDSW. Although both methods show close performance in HOTA and MOTA with VirConv [19], CasA [20], and PointRCNN [22], CasTrack [1] performance drops dramatically in PV-RCNN [23] with a margin of more than $6\%$ in MOTA and $4\%$ in HOTA.
Our proposed method, while maintaining the lowest IDSW compared to CasTrack [1], consistently outperforms across all detectors. These results validate the contribution of the proposed tracking approach, demonstrating its exceptional performance across various detectors. Additionally, the proposed approach achieves higher frame per second (fps) than CasTrack [1] across the detectors, with an fps enhancement reaching $+59\%$ with the Second [21] detector.
Another evaluation is conducted on the WOD testing dataset to measure the generalization of the approach. The dataset consists of simple to moderate tracking conditions labeled “Level 1" and challenging conditions labeled “Level 2" in Table 2. Our proposed solution shows superior tracking performance on the recent benchmark, including CasTrack [1], with a reduction in the false positives ratio (FP).
@@
Table 2: Tracking performance on Waymo Open dataset testing split with the latest MOT benchmarks. Methods that use the CasA [20] detector are marked by (*)
@@
CasTrack* [1] 66.95% 7.74% 63.66% 7.36%
ImmotralTrack [34] 63.77% 8.94% 60.55% 8.48%
SimpleTrack [35] 63.53% 8.66% 60.30% 8.75%
CenterPoint [36] 62.58% 9.03% 59.38% 9.41%
AlphaTrack [37] 58.86% 10.10% 55.67% 9.55%
RobMOT* (Ours) 67.02% 7.60% 63.67% 7.22%
In conclusion, the proposed approach excels in tracking stability and accuracy, significantly reducing ID switches and enhancing fps performance across various detectors. It outperforms recent state-of-the-art methods like CasTrack [1] on both the KITTI and WOD datasets, demonstrating robust generalization and efficiency. The quantitative results validate the contribution and exceptional performance of the proposed framework.

### v4

@@ 3.2 Trajectory Association and State Estimation with a Refined Kalman Filter for Noise Mitigation
It has been observed that state estimation drifts for objects encountering occlusion or leaving the sensor field of view (off-scene). This phenomenon is also observed in the literature, as displayed in Figure 1 when the estimated trajectory of a parked car with ID 4 drifts away from the expected location during the occlusion. Figure 3 shows another scenario when two parked cars with ID 0 and 2 drift from the actual position after being off-scene. Our investigation of the phenomenon concludes that this deviation is caused by trembling in detection bounding boxes around objects’ locations, which results in illusion movement for these objects that eventually impacts state update and prediction in KF.
This noise is modeled using the algebraic deviation, shown in Equation 1, between the detected bounding boxes $z^{det}_{ij}=(x_{ij},y_{ij})$ and the actual location of the objects $z^{gt}_{ij}=(x_{ij},y_{ij})$ across $k$ observations for $N$ objects using KITTI training dataset.
@@
As depicted in Figure 4, the error distribution shows a Gaussian shape for the noise of the VirConv detector [2] across $x$ and $y$ LiDAR coordinates, which facilitates its integration into KF. Consequently, $2\times 2$ covariance matrix $D_{t}$ is driven that represents the impact of noise on the detections (measurements). In particular, noise varies from one detector to another as to the variance in detection performance; hence, $D_{t}$ should be recomputed as the employed detector changes. By applying Equation 1 on $y$ coordinate, a covariance matrix of the noise is obtained, Equation 2.
@@
Since this noise is attached to the incoming detections $z_{t}$ at time $t$, it harms the computation of the residual term $\hat{y}_{t}$, which is the difference between the current measurement $z_{t}$ and state estimation $\hat{x}_{t|t-1}$ predicted at $t-1$. Consequently, the noise impact on the residual term leads to incorrect updates for the motion parameters of the object. The innovation covariance $S_{t}$ is responsible for weighting the innovation residual $\hat{y}_{t}$ through the Kalman gain $K_{t}$. Therefore, the covariance matrix $D_{t}$ needs to be added to the innovation covariance $S_{t}$ to compensate for the impact of noise on the updated state. Hence, the updated state estimation $\hat{x}_{t|t}$ will consider the impact of the deviation noise on the observation.
@@
The main reason for not using $R_{t}$ instead is because it models the incorrectness of the measurement for the point cloud from LiDAR with respect to the real world, while $D_{t}$ models the incorrectness of the detection process with respect to the LiDAR point cloud, which means they are independent. Equation 3 demonstrates the updated state of KF after adapting $D_{t}$ noise. $H_{t}$ is the observation model to map from estimation to measurement space at time $t$, while $P_{t|t-1}$ is the estimation uncertainty at time $t$ given measurement at $t-1$. Section 3.4 will describe how $P_{t|t}$ contribute to trajectory termination in the cache.
3.3 Trajectory Validity Mechanism
The objective is to monitor the validity (legitimacy) of the existing trajectories in the system while identifying invalid ones, called ghost tracks. Each track has a certainty score that represents its legitimacy. This score is updated whenever a new observation of the track is available. The certainty score changes based on the likelihood of the trajectory being a ghost track. This mechanism is inspired by humans’ ability to perceive and identify unfamiliar objects. With multiple observations, human perception can verify and identify unfamiliar objects. By observing the characteristics of ghost tracks, the ghost tracks are usually formed from intermittent sequences of observations with low to moderate detection scores. Based on these characteristics, the certainty score for track $i$ changes when a new observation is associated with the track at time $t$. The certainty score changes according to the detection score $s_{i}^{t}$ and the duration of absence $d^{i}_{t}$ since the last observation associated with the track.
@@ 3.3.2 Absence duration
The penalty term $\frac{d^{i}_{t}}{s^{i}_{t}}$ lowers the certainty score as the absence duration $d^{i}_{t}$ of a new observation for the track increases. This term will diminish the impact of the reward with an extended absence duration that indicates a high potential of being a ghost trajectory. However, this term can harm legitimate tracks that encounter occlusion. Since observations of legitimate tracks tend to have high detection scores $s^{i}_{t}$, the inverse proportion is included in the detection score to decrease the impact of this term on the legitimate tracks.
@@ 3.3.3 Temporal validity
A visual representation of the distribution of Equation 4 is highlighted in Figure 5; the relation between the detection score, the absence duration, and its impact on the certainty score. Even with low detection scores, Equation 4 returns a higher reward without decay ($e^{-}{d_{t}^{i}}$) or penalty ($\frac{d_{t}^{i}}{s_{t}^{i}}$) as long as the observation is consistent (too small to zero absence duration). Thus, the trajectory of true positive detections with low detection scores caused by partial occlusion or distance can still be verified.
This mechanism monitors the validity of trajectories in the system individually until a certainty score of a trajectory exceeds a pre-defined threshold $\alpha_{legit}$ that indicates the trajectory is legitimate. In this case, the validation mechanism changes the trajectory status from “not confirmed” to “confirmed” and terminates the validation mechanism to this trajectory. Thus, confirmed trajectories cannot be changed to “not confirmed" as the mechanism is terminated for these tracks.
@@ 3.4 Trajectory Termination and Caching
Figure 6: An illustration of spatial state estimation uncertainty for a car in a 2D dimension shows high estimation uncertainty along the x-axis while low uncertainty on the y-axis.
The cache holds all active trajectories. In the literature, a trajectory is terminated or discarded when there are no further observations for that trajectory for several frames. Generally, it is not trivial to define an absence threshold, as it could risk eliminating objects under prolonged occlusion or maintaining ghost tracks for a long time. Instead, trajectories are discarded when there is high uncertainty about their estimation, as that uncertain estimation is more likely to be mismatched. Each trajectory has an estimation uncertainty $P_{t|t}$ from Equation 3 on $x$ and $y$ direction that forms an ellipse, as shown in Figure 6. The trajectory is discarded when estimation uncertainty ${diag(P_{t})_{x},diag(P_{t})_{y}}$ becomes high in either direction, exceeding $\sigma_{est\_certainity}=4$ meters. This strategy serves trajectories under occlusion and fast elimination for ghost tracks. As per our observation, ghost tracks tend to have high growth in estimation uncertainty because of their limited observations, while legitimate tracks maintain low estimation uncertainty, even those under occlusion.
@@ 4 Results and Discussion
The framework is evaluated using two benchmark datasets for autonomous cars: KITTI [5] and Waymo Open Dataset [30] (WOD). The KITTI dataset consists of $50$ sequences, split into $21$ training sequences and $29$ testing sequences recorded around Karlsruhe, Germany, that involve a variety of challenging conditions. Meanwhile, WOD contains about $1150$ segments recorded in various locations across the United States.
@@
4.1 Implementation Details
This research is implemented in C++17. The modules have been implemented from scratch to ensure the efficiency and precise implementation of the proposed methods. The Eigne3 library is used to perform the algebraic procedures discussed in this work. Meanwhile, OpenCV 4.6 and PCL 1.14 are used to visualize the result. The tracking evaluation of KITTI validation and training datasets was done using their official evaluation tool [31]. The experiments are conducted on a machine with a processor AMD Ryzen 9 and 32 GB RAM; no GPU is involved since the framework utilizes a single CPU.
@@ 4.2 Evaluation with Benchmarks
The primary metrics used in the evaluation are HOTA and MOTA (See Equation 5). MOTA metric relies mainly on the false positives (FP) (ghost tracks), false negatives (FN) (missed tracks), and switch in tracks’ identity (IDSW). Since a single detector is unified among experiments, FN remains the same as long as the detector is not changed so that the enhancement in MOTA comes from fewer ghost tracks (FP) and IDSW. On the other hand, HOTA is concerned about detection accuracy (DetA) and association accuracy (AssA). AssA represents the average localization of the state estimation and the actual object location based on how well the state estimation bounding box overlaps with the actual object location ($IoU_{matched}$). Accordingly, the improvement in HOTA while unifying the detector indicates improvement in the state estimation localization.
$\begin{split}\text{MOTA}&=1-\frac{\sum_{t}(FN_{t}+FP_{t}+IDSW_{t})}{\sum_{t}GT_{t}}\\ \text{HOTA}&=\sqrt{\text{DetA}\times\text{AssA}}\\ \text{AssA}&=\frac{\sum\text{IoU}_{\text{matched}}}{\text{Total correct matches}}\end{split}$ (5)
To this end, MOTA can measure the contribution of the proposed trajectory validity mechanism for ghost tracks. At the same time, HOTA can indicate the impact of the proposed noise mitigation term $D_{t}$ on enhancing state estimation localization for objects in the KF.
@@ 4.2.1 Quantitative Evaluation
Table 1: The table compares the latest benchmarks on the testing split in the KITTI leaderboard. The highest performance among classical methods at each metric is highlighted in red, the second highest in blue, the third highest in green, and the fourth highest in violet. Methods that use the CasA [4] detector are marked by (*), while ($\dot{+}$) is used for methods that use the VirConv [2] detector.
@@
Method HOTA MOTA AssA AssRe IDSW TripletTrack [32] 73.58% 84.32% 74.66% 77.3% 322 PolarMOT [33] 75.2% 85.1% 76.95% 80.0% 462 DeepFusion-MOT [6] 75.5% 84.6% 80.1% 82.6% 84 Mono-3D-KF [34] 75.5% 88.5% 77.6% 80.2% 162 PC3T [35] 77.8% 88.8% 81.6% 84.8% 225 MSA-MOT [36] 78.5% 88.0% 82.6% 85.2% 91 UG3DMOT* [37] 78.6% 87.98% 82.28% 85.36% 30 CasTrack* [1] 77.3% 86.29% 80.29% 83.12% 184 $\dot{+}$ [3, 1] 79.9% 89.1% 82.6% 85.6% 201 PC-TCNN [38] 80.9% 91.7% 84.1% 87.5% 37 RobMOT* (Ours) 81.22% 90.5% 85.8% 89.7% 7 $\dot{+}$ (Ours) 81.76% 91.02% 85.6% 89.25% 7
The proposed method is evaluated quantitatively on the KITTI testing dataset with the latest published state-of-the-art tracking methods listed in the KITTI leaderboard [5]. Since the proposed method is an online method (real-time) while the leaderboard combines online and offline methods, Table 1 reports the online methods (classical and deep learning methods). The framework outperforms recent online benchmarks in almost all metrics, including the recent deep learning approach PC-TCNN [38], using either CasA [4] or Virconv [2] detector. As our approach is a classical method, the evaluation will concentrate on the recent classical benchmark method [3, 1] (VirConvTrack/CasTrack) because they achieve the highest performance as per Table 1 and the KITTI leaderboard [5].
RobMOT outperforms CasTrack and VirConvTrack in all metrics using the same detectors, with a margin reaching $3.92\%$ in HOTA, $5.79\%$ in AssA, and $6.85\%$ in association recall (AssRe) with the CasA detector. Notably, the proposed method shows a significant reduction in IDSW in both detectors, which shows its tracking stability.
A more intensive evaluation study is conducted on the KITTI evaluation dataset across five different detectors with the method [1, 3] as it achieves the highest tracking performance among the classical methods. Figure 7 represents the outcome of this study in HOTA, MOTA, and IDSW metrics. The proposed method consistently outperforms the benchmark [1, 3] in all the metrics across the detectors. The performance margin in HOTA reaches $8.7\%$ and $16.1\%$ in MOTA with PV-RCNN detector [29]. Moreover, our method shows a significant reduction in IDSW, as shown in Figure 7 (the third graph), which supports results in Table 1 on the test dataset. This significant performance drop in work [1] shows the dependency of the method on the detector performance.
@@
Table 2: Tracking performance on Waymo Open Dataset testing split with the latest online MOT benchmarks. The color code used in the table to highlight the first-highest (red) and second-highest (green) performance for benchmarks use the same detector. Benchmarks with the CasA [4] detector are marked by (*) while the CTRL [39] detector is marked by (**).
@@
SimpleTrack [40] $[0,30)$ 86.23% 0.02% 85.42% 0.02%
ImmotralTrack [41] $[0,30)$ 86.35% 0.00% 85.55% 0.00%
FastPoly [42]* $[0,30)$ 87.89% 0.01% 87.07% 0.01%
CasTrack [1]* $[0,30)$ 88.05% 0.02% 87.24% 0.02%
CasTrack [1]** $[0,30)$ 91.98% 0.01% 91.37% 0.01%
RobMOT* (Ours) $[0,30)$ 88.10%(+0.05%) 0.01%(–0.01%) 87.29%(+0.05%) 0.01%(-0.01%)
RobMOT** (Ours) $[0,30)$ 92.17%(+0.19%) 0.00%(-0.01%) 91.59%(+0.22%) 0.00%(-0.01%)
SimpleTrack [40] $[30,50)$ 64.09% 0.08% 61.14% 0.08%
ImmotralTrack [41] $[30,50)$ 64.41% 0.01% 61.45% 0.01%
FastPoly [42]* $[30,50)$ 66.72% 0.05% 63.46% 0.05%
CasTrack [1]* $[30,50)$ 67.42% 0.06% 64.38% 0.06%
CasTrack [1]** $[30,50)$ 78.07% 0.08% 75.08% 0.08%
RobMOT* (Ours) $[30,50)$ 67.72%(+0.30%) 0.02%(+0.04%) 64.63%(+0.25%) 0.02%(-0.04%)
RobMOT** (Ours) $[30,50)$ 78.78%(+0.71%) 0.03%(-0.05%) 75.99%(+0.91%) 0.02%(-0.06%)
SimpleTrack [40] $[50,+inf)$ 36.28% 0.14% 32.71% 0.12%
ImmotralTrack [41] $[50,+inf)$ 36.59% 0.02% 33.02% 0.02%
FastPoly [42]* $[50,+inf)$ 40.99% 0.10% 37.18% 0.09%
CasTrack [1]* $[50,+inf)$ 41.81% 0.08% 37.90% 0.07%
CasTrack [1]** $[50,+inf)$ 58.47% 0.14% 53.37% 0.12%
RobMOT* (Ours) $[50,+inf)$ 42.39%(+0.58%) 0.03%(-0.05%) 38.37%(+0.47%) 0.03%(-0.04%)
RobMOT** (Ours) $[50,+inf)$ 59.74% (+1.27%) 0.04%(-0.10%) 55.14%(+1.77%) 0.03%(-0.09%)
Another evaluation is conducted on the WOD testing dataset to measure the generalization of the approach. The dataset consists of simple to moderate tracking conditions labeled “Level 1" and challenging conditions labeled “Level 2" in Table 2. Table 2 represents the tracking performance of the state-of-the-art methods on WOD using MOTA and Mismatch (miss association) metrics. The table consists of tracking evaluations for objects between different distance ranges from the AV: 0 to 30 meters, 30 to 50 meters, and more than 50 meters.
The proposed method, RobMOT, achieves the highest tracking performance in both level 1 and level 2 in MOTA and Mismatch. The performance margin with the other benchmarks increases as the tracking distance range increases, reaching $1.77\%$ in MOTA for objects in the range $[50,+inf)$. That shows the method’s capability to track distant objects over the recent benchmarks. Moreover, the proposed method consistently outperforms CasTrack [1] while unifying the employed detectors (CasA [1] and CTRL [39]) for both methods.
In conclusion, the proposed approach excels in tracking stability and accuracy, significantly reducing ID switches. It outperforms recent state-of-the-art methods [1, 3] on both the KITTI and WOD datasets, demonstrating robust generalization and efficiency. The quantitative results validate the contribution and exceptional performance of the proposed framework.